OpenStudy (anonymous):

Find f if f′′(x)=sinx+cosx, f′(0)=4, and f(0)=5 find f(x) ------------------------------ I got the integral and (double?) integral: -cosx+sinx and -sinx-cosx respectivly. I have no cluee how to solve the rest here. I got it with another similar problem but how do i do it with 2 integrals?

OpenStudy (misty1212):

you are missing the constants which you find like before

OpenStudy (misty1212):

\[f'(x)=-\cos(x)+\sin(x)+C\] and you need \(C\) via \[f'(0)=4\]

OpenStudy (anonymous):

hey misty sorry im not so good at these lol so i would need to to plug (0) into x and set it equal to 4 right?

OpenStudy (misty1212):

yes, just like the other one no need to be sorry though

OpenStudy (anonymous):

so i would have 5=C correct? for the first part?

OpenStudy (anonymous):

would this mean I would have -sinx-cosx+5x=f(x)?

OpenStudy (misty1212):

\[4=-1+C\\ C=5\] ok good

OpenStudy (misty1212):

and yes but don't forget the +C you have to do this dog an pony show again

OpenStudy (misty1212):

\[f(x)=-\cos(x)-\sin(x)+5x+C\] etc etc

OpenStudy (misty1212):

i have a question

OpenStudy (misty1212):

why don't you use this one?

OpenStudy (anonymous):

lol I think I will

OpenStudy (anonymous):

but one more question

OpenStudy (anonymous):

So I ended up with 6=C for f(x) but I feel that is eerily wrong

OpenStudy (misty1212):

lets check

OpenStudy (misty1212):

\[f(x)=-\cos(x)-\sin(x)+5x+C\] \[f(0)=-1+C=5\\ C=6\] what is wrong with that ?q

OpenStudy (anonymous):

...Nothing? i think?